### v1

# Measurable Krylov Spaces and Eigenenergy Count in Quantum State Dynamics

Abstract
We define a new Krylov space for the time evolution of a quantum state, which is both quantum-mechanically measurable. Based on our defined spaces, an effective dimension is introduced that provides insight into how expressive the quantum system is. The understanding of the time evolution operator as a map onto a Krylov space is evident from the Taylor expansion of said operator. Current literature focuses on computing different powers of the Hamiltonian to construct a basis for the Krylov state space. One challenge this approach faces is that computing higher powers of the Hamiltonian becomes increasingly difficult for larger systems. Since each basis state is constructed from a different power of the Hamiltonian, the basis states have varying lengths, making the ortho-normalization process challenging. Furthermore, computing high powers of the Hamiltonian increases rounding errors, which can lead to a wrongly increased Krylov space dimension. The first part of our work challenges these issues by defining an equivalent space, where the original basis consists of states of the same length. We demonstrate that a set of different time-evolved states can be used to construct a basis. We subsequently verify the results through numerical analysis, demonstrating that every time-evolved state can be reconstructed using the defined vector space. Based on this new space, we define an effective dimension and analyze its influence on finite-dimensional systems. Our method gives insight into the basis construction of the time evolution operator, which can be computed experimentally. We further show that the Krylov space dimension is equal to the number of pairwise distinct eigenvalues of the Hamiltonian, enabling a method to determine the number of eigenenergies the system has experimentally.
1 Introduction
Our understanding of quantum systems primarily arises from the Schrödinger equation for pure states and the von-Neumann equation for mixed states. The comprehension of how quantum states evolve over time is crucial for various technologies. Therefore, recent years have seen increased attention towards understanding the space in which quantum state evolution occurs.
In this work, we aim to define a measurable space for quantum time evolution under the Schrödinger equation. Our research explores operator complexity, as initially introduced in PAR19 , and spread complexity, first explored in BAL22 . A unified perspective on state and operator complexities was later presented in ALI23 , where the authors demonstrated that the time evolution of observables or states can be viewed as a mapping onto a Krylov space.
Research in the understanding of Krylov complexity has been done in regards to finite and infinite dimensional quantum systems, field theories, quantum chaos, open quantum systems, entanglement, and more (see BAR19 ; BHA23 ; CAM23 ; CAP21 ; DYM20 ; DYM21 ; IIZ23 ; BAL22 ; MUC22 ; GUO22A ; PAT22 ; FAN22 ; HEV22 ; RAB22 ; RAB22 ; RAB22a ; NIZ23 ; LIU23b ). Spread complexity has been analyzed concerning Nielsen complexity, describing topological phases, chaotic dynamics, finite and infinite dimensional quantum systems, and so forth BAL22 ; BAL22a ; BAL23 ; CRA23 ; AGU24 ; CAP22 ; CAP22a ; ERD23 ; GIL23 ; NAN24 ; PAL23 ; BHA22 ; CHA23 ; GAU23 ; NIE06a ; BHA24 .
The numerical computation of these complexity measures makes use of the Lanczos algorithm, which is prone to rounding error. Another problem this approach has is that the Krylov space has to be constructed through classical simulation of the system. With increasing qubit sizes or for large quantum systems, this becomes infeasible. Both problems will be addressed in this work, as we introduce a space that is equivalent to the Krylov space discussed in BAL22 . We further show that the basis of our space can be computed through quantum mechanical measurements and that the dimension is equal to the number of pairwise distinct eigenvalues of the Hamiltonian. This gains prominence as quantum computers attain sizes beyond the computational capabilities of classical machines.
Quantum computation tries to make use of the exponential scaling in space dimension for computing resources NIE11 . The most prominent quantum algorithm is Shor’s algorithm, which can perform prime factorization more efficiently than a classical computer SHO94 . The simulation of quantum systems on a quantum computer, which are classically difficult to simulate, appears increasingly plausible with the continuous growth in qubit sizes from 127 qubits in 2021 to 1121 qubits in 2023 on IBMs superconducting devices. Another important aspect of classical computation is machine learning, and therefore, the question arises: how can machine learning algorithms be implemented using quantum systems?
One approach to qunatum machine learning is variational quantum machine learning, where a quantum network is trained on a quantum computer. The second scheme is quantum reservoir computing, which tries to utilize the large Hilbert space as a computing resource for time series forecasting. This scheme is especially challenging, because of the collapse of the state after measurement. Some solutions to this problem have been proposed in CIN24 ; MUJ23 . Lastly quantum extreme learning machines use quantum systems for computation of finite data. Several experiments in quantum machine learning have been conducted on quantum hardware, as detailed in LI15 ; SAG21 . The comparative analysis of machine learning networks remains an active field of research, and the question of expressivity in Quantum Machine Learning (QML) remains unresolved. The central inquiry revolves around the methodology for comparing two Hamiltonians or two quantum machine learning networks in terms of the dimension of the space that comprehensively spans all future states.
Section 2 of this work will introduce the spread complexity shortly, which was proposed in BAL22 . The development of a measurable Krylov space will be discussed in Section 3. We will prove our main theorems and construct a Krylov space that is computable through quantum mechanical measurements (Theo. 2 and Theo. 3). Afterwards, we will show that the dimension of this space is equivalent to the number of pairwise distinct eigenvalues of the Hamiltonian in Theo. 4. The results will be first verified by the reconstruction capability of the defined space through numerical analysis, see Section 4.
2 Fundamentals of spread complexity
This chapter will start with Def. 1, which defines the time-evolution and the function of interest. The observation of the time-evolution operator being a map onto a Krylov space was discussed in PAR19 ; BAL22 . Here, the proof for Theo. 1 was given. Building upon this, the Lanczos algorithm is performed, and different complexity measures are introduced for various systems LAN50 ; PAR19 ; BAL22 .
The Schrödinger equation for the time-independent Hamiltonian $H$ with initial condition |$\ket{\Psi(0)}$ is given by:
$\partial_{t}\ket{\Psi(t)}=-iH\ket{\Psi(t)}$
|$\ket{\Psi(0)}:=\ket{\Psi_{0}}$ (1)
The solution to this equation is
|$\ket{\Psi(t)}=e^{-iHt}\ket{\Psi_{0}}$ (2)
and the density matrix $\rho\in\mathbb{C}^{N\times N}$ is constructed by
$\rho=\ket{\Psi}\bra{\Psi}.$ | . (3)
A measurement in quantum mechanics is the application of a Hermitian matrix $O$, also called an observable, with the eigenvalue equation $O\ket{o_{i}}=o_{i}\ket{o_{i}}$ onto the system. Only the eigenvalues of $O$ can be measured, and which eigenvalue is measured is given by the probability
$P(m=o_{i})=|\bra{o_{i}}\ket{\Psi}|^{2}.$ (4)
The expectation value of the observable $O$ in regard to the state |$\ket{\Psi}$ is given by
$\expectationvalue{O}=\bra{\Psi}O\ket{\Psi}=\mathrm{Tr}(\rho O).$ (5)
This work will construct a basis |$\ket{w_{1}},\ket{w_{2}},..\ket{w_{m}}$ for some Hamiltonian $H$ and starting state |$\ket{\Psi_{0}}$, such that
|$\ket{\Psi(t)}\in\mathrm{Span}(\ket{w_{0}},\ket{w_{1}},..,\ket{w_{m}})$ (6)
holds for all $t$. The established basis should also be attainable through quantum mechanical measurements, particularly for systems where the precise Hamiltonian is unknown or when the system’s size surpasses classical simulation capabilities. Quantum mechanics commonly employs the representation of complex vectors within the Hilbert space, denoted as ${\mathcal{H}=\mathbb{C}^{N}}$. We will specifically focus on the quantum-mechanical state throughout our discussion.
Definition 1 (Time Evolution).
Let $t\in\mathbb{R}_{\geq 0}$, |$\ket{\Psi_{0}},\ket{\Psi(t)}\in\mathbb{C}^{N}$ and $H\in\mathbb{C}^{N\times N}$ be an Hermitian matrix, meaning $H=H^{\dagger}$. The time evolution $F:\mathbb{R}\times\mathbb{C}^{N}\rightarrow\mathbb{C}^{N}$ with initial condition |$\ket{\Psi(t=0)}=\ket{\Psi_{0}}$ is defined by
|$\ket{\Psi(t)}=F(\ket{\Psi_{0}})=e^{-iHt}\ket{\Psi_{0}}.$ (7)
Def. 2 gives the general definition of a Krylov space and how to construct this space.
Definition 2 (Krylov Space).
Let $f:\mathbb{C}^{N}\rightarrow\mathbb{C}^{N}$ be a linear function, i.e., for all $\mathbf{v},\mathbf{w}\in\mathbb{C}^{N}$ and $\alpha\in\mathbb{C}$, the following holds:
$f(\mathbf{v}+\mathbf{w})$ $=f(\mathbf{v})+f(\mathbf{w}),\quad f(\alpha\mathbf{w})=\alpha f(\mathbf{w})$ (8)
Let $\mathbf{v}_{0}\in\mathbb{C}^{N}\setminus\{\mathbf{0}\}$ be a vector. Then there exists a uniquely defined smallest $m\leq N$ such that the vectors
$\mathbf{v}_{0},f(\mathbf{v}_{0}),f^{2}(\mathbf{v}_{0}),\ldots,f^{m-1}(\mathbf{v}_{0})$
are linearly independent, and the vectors
$\mathbf{v}_{0},f(\mathbf{v}_{0}),f^{2}(\mathbf{v}_{0}),\ldots,f^{m}(\mathbf{v}_{0})$ (9)
are linearly dependent. The number $m$ is called the grade of the vector $\mathbf{\mathbf{v}_{0}}$ with respect to $f$. The letter $m$ will always be used to denote a grade. The $j$-th Krylov space of $f$ with respect to $\mathbf{v}_{0}$ is defined as
$\mathrm{K}_{j}(f,\mathbf{v}_{0})=\mathrm{Span}\{\mathbf{v}_{0},f(\mathbf{v}_{0}),\ldots,f^{j-1}(\mathbf{v}_{0})\}\subseteq\mathbb{C}^{N}.$ (10)
Note that $\mathrm{K}_{M}(f,\mathbf{\mathbf{v}_{0}})=\mathrm{K}_{m}(f,\mathbf{\mathbf{v}_{0}})$ holds for all $M\geq m$. If the chosen initial state $\mathbf{v}_{0}$ and function $f$ are clear, the corresponding Krylov space with grade $m$ will be denoted as $\mathrm{K}_{m}$.
Theo. 1 will give the proof that the time-evolution, as defined in Def. 1, maps the initial state |$\ket{\Psi_{0}}$ onto a Krylov space, as discussed in PAR19 ; BAL22 .
Theorem 1.
[Time Evolution as a Map onto a Krylov Space] Assume the system from Def. 1. Let $t\in\mathbb{R}$ and let |$\ket{\Psi(t)}=\exp(-iHt)\ket{\Psi_{0}})$ be the time-evolved state by $t$, where |$\ket{\Psi_{0}}\neq 0$. For any $t\in\mathbb{R}$
|$\ket{\Psi(t)}\in\mathrm{K}_{m}(-iH,\ket{\Psi_{0}})=\mathrm{Span}\{f^{0}(\ket{\Psi_{0}}),\ldots,f^{m-1}(\ket{\Psi_{0}})\}$
holds true, where $m\in\mathbb{N}$ is the grade of the starting state |$\ket{\Psi_{0}}$ in regards to the linear function $f:=-iH$PAR19 ; BAL22 .
Proof.
We know that |$\ket{\Psi(t)}=\exp(-iHt)\ket{\Psi_{0}}$ is given. With the introduction of the linear function $f:\mathbb{C}^{N}\rightarrow\mathbb{C}^{N}$ as $f(\ket{\Psi_{0}}):=-iH\ket{\Psi_{0}}$, the Taylor expansion of the time evolution can be rewritten as
|$\ket{\Psi(t)}$ $=F(\ket{\Psi_{0}})=e^{-iHt}\ket{\Psi_{0}}$
$=\sum_{k=0}^{\infty}(-iH)^{k}\frac{t^{k}}{k!}\ket{\Psi_{0}}=\sum_{k=0}^{\infty}f^{k}(\ket{\Psi_{0}})\frac{t^{k}}{k!}.$ (11)
|$\ket{\Psi(t)}$ is a superposition of vectors $f^{k}(\ket{\Psi_{0}})$ with amplitudes $\frac{t^{k}}{k!}$, such that
|$\ket{\Psi(t)}\in\mathrm{Span}\{f^{0}(\ket{\Psi_{0}}),f^{1}(\ket{\Psi_{0}})t,f^{2}(\ket{\Psi_{0}})\frac{t^{2}}{2!},\ldots\}$ (12)
holds. In the trivial case with $t=0$, |$\ket{\Psi(t)}=\ket{\Psi_{0}}$ follows. In the case of $t>0$, the values $\frac{t^{k}}{k!}>0$, such that the span can be represented as
$\mathrm{Span}(f,\ket{\Psi_{0}})$ $=\mathrm{Span}\{f^{0}(\ket{\Psi_{0}}),f^{1}(\ket{\Psi_{0}}),f^{2}(\ket{\Psi_{0}}),\ldots\}.$ (13)
By Def. 2 and the linear property of $f$, there exists an $m\leq N$, such that
|$\ket{\Psi_{0}},f(\ket{\Psi_{0}}),f^{2}(\ket{\Psi_{0}}),\ldots,f^{m-1}(\ket{\Psi_{0}})$ (14)
are linearly independent, and such that
|$\ket{\Psi_{0}},f(\ket{\Psi_{0}}),f^{2}(\ket{\Psi_{0}}),\ldots,f^{m}(\ket{\Psi_{0}})$ (15)
are linearly dependent. This implies that
$\mathrm{Span}(f,\ket{\Psi_{0}})$ $=\mathrm{Span}\{f^{0}(\ket{\Psi_{0}}),f^{1}(\ket{\Psi_{0}})t,f^{2}(\ket{\Psi_{0}})\frac{t^{2}}{2!},\ldots\}$
$=\mathrm{Span}\{f^{0}(\ket{\Psi_{0}}),f^{1}(\ket{\Psi_{0}}),f^{2}(\ket{\Psi_{0}}),\ldots\}$
$=\mathrm{Span}\{f^{0}(\ket{\Psi_{0}}),f^{1}(\ket{\Psi_{0}}),\ldots,f^{m-1}(\ket{\Psi_{0}})\}$
$=\mathrm{K}_{m}(-iH,\ket{\Psi_{0}}),$ (16)
and therefore |$\ket{\Psi(t)}\in\mathrm{K}_{m}(-iH,\ket{\Psi_{0}})$, where $m$ is the grade of the vector |$\ket{\Psi_{0}}$ concerning $f:=-iH$. ∎
In current analysis of quantum complexity the above defined spaces $\mathrm{K_{m}}$ are used to define a complexity measure BAL22 . First, the Lanczos algorithm is performed on this space to construct the space $\mathrm{K_{m}}=\mathrm{Span}(\ket{k_{1}},\ket{k_{2}},..,\ket{k_{m}})$. The state representation of any time evolved state |$\ket{\Psi(t)}$ is given by
|$\ket{\Psi(t)}=\sum_{n=0}^{m}\alpha_{n}(t)\ket{k_{n}}.$ (17)
The spread complexity $C_{S}$ is am measure of the spread of the state over the basis consisting of different powers of the Hamiltonian and is defined as
$C_{S}(t)=\sum_{n=1}^{m}n\absolutevalue{\alpha_{n}(t)}^{2}.$ (18)
3 Measurable Krylov spaces
In our work, we propose different basis for the Krylov space, given in Theo. 2 and LABEL:theorem:2-2_main_proof. We show that instead of constructing the Krylov basis as in Theo. 1, the space can be constructed from different time-evolved states of the system. Lemma 1 shows that the constructed space is invariant under a global phase, thereby becoming quantum mechanically measurable. With Lemma 2, we discuss how different evolution times can be chosen to construct the space discussed in Theo. 2 and Theo. 3. To verify the mathematical results, we construct another basis, which implies that the number of pairwise distinct eigenvalues should coincide with the dimension of the Krylov space (see Theo. 4).
Definition 3 (Samplings of Krylov Space).
Assume |$\ket{\Psi_{0}}\in\mathbb{C}^{N}$, |$\ket{\Psi_{0}}\neq 0$, and $H\in\mathbb{C}^{N\times N}$, and consider the time evolution for $T>0$ such that
|$\ket{\Psi(t)}=e^{-iHT}\ket{\Psi_{0}}$ (19)
holds. Let $\mathrm{K}_{m}(-iH,\ket{\Psi_{0}})$ be the Krylov space of $f:=-iH$ with respect to |$\ket{\Psi_{0}}$, where $m$ is the grade of |$\ket{\Psi_{0}}$ concerning $f:=-iH$. With discrete times $0=t_{0}<t_{1}<\ldots<t_{M-1}=T$, the vectors
|$\ket{g_{i}}=\ket{\Psi(t_{i})}=e^{-iHt_{i}}\ket{\Psi_{0}}\mathrm{,~{}for~{}}i=1,\ldots,M,$ (20)
are introduced, addressed as the sampling vectors of the time evolution.
Theorem 2 (Approximation of Krylov Space).
Let $0=t_{0}<t_{1}<\ldots<t_{n-1}<T_{P}$ be some times. Next, the following vectors are introduced:
${h_{i}^{(n)}}(\ket{\Psi_{0}})$ $:=\sum_{j=0}^{n-1}{f^{j}(\ket{\Psi_{0}})}\frac{t_{i}^{j}}{j!}$ (21)
$\mathrm{with~{}}f^{j}(\ket{\Psi_{0}})$ $:=(-iH)^{j}\ket{\Psi_{0}}.$
The span of the vectors ${h_{i}^{(n)}}(\ket{\Psi_{0}})$ is given by
$\mathrm{H}_{n}^{n}$ $=\mathrm{Span}({h_{0}^{(n)}}(\ket{\Psi_{0}}),{h_{1}^{(n)}}(\ket{\Psi_{0}}),\ldots,{h_{n-1}^{(n)}}(\ket{\Psi_{0}})).$ (22)
Then for all $n\in\mathbb{N}$, it holds that
$\mathrm{H}_{n}^{n}=\mathrm{K}_{n}.$ (23)
Proof.
Assume $n\in\mathbb{N}$. Then the vectors ${h_{i}^{(n)}}(\ket{\Psi_{0}})$ can be represented with the vectors ${f^{j}(\ket{\Psi_{0}})}$ as
${h_{i}^{(n)}}(\ket{\Psi_{0}})=\sum_{j=0}^{n-1}{f^{j}(\ket{\Psi_{0}})}\frac{t_{i}^{j}}{j!}$
$=({f^{0}(\ket{\Psi_{0}})},{f^{1}(\ket{\Psi_{0}})},\ldots,{f^{n-1}(\ket{\Psi_{0}})})\begin{pmatrix}1\\ t_{i}/1!\\ t_{i}^{2}/2!\\ :\\ t_{i}^{n-1}/({n-1})!\end{pmatrix}.$ (24)
Writing the vectors as column vectors in a matrix results in
$({h_{0}^{(n)}}(\ket{\Psi_{0}}),\ldots,{h_{n-1}^{(n)}}(\ket{\Psi_{0}}))$
$=({f^{0}(\ket{\Psi_{0}})},{f^{1}(\ket{\Psi_{0}})},\ldots,{f^{n-1}(\ket{\Psi_{0}})})\Theta$ (25)
with
$\Theta$ $=\begin{pmatrix}1&1&\ldots&1\\ t_{1}/1!&t_{2}/1!&\ldots&t_{n-1}/1!\\ t_{1}^{2}/2!&t_{2}^{2}/2!&\ldots&t_{n-1}^{2}/2!\\ :&:&:&:\\ t_{1}^{n-1}/(n-1)!&t_{2}^{n-1}/(n-1)!&\ldots&t_{n-1}^{n-1}/(n-1)!\end{pmatrix}.$ (26)
For $i\neq j$, $t_{i}\neq t_{j}$ is given. It follows that all column vectors of $\Theta$, called Vandermonde matrix, are linearly independent, and therefore $\Theta$ is invertible. The vectors of $\mathrm{K}_{m}$ can then be represented with the inverse $\Theta^{-1}$ as
$({f^{0}(\ket{\Psi_{0}})},\ldots,{f^{n-1}(\ket{\Psi_{0}})})$ $=({h_{0}}(\ket{\Psi_{0}}),\ldots,{h_{n-1}}(\ket{\Psi_{0}}))\Theta^{-1}.$ (27)
As all vectors ${f^{0}(\ket{\Psi_{0}})},\ldots,{f^{n-1}(\ket{\Psi_{0}})}$ of $\mathrm{K}_{n}$ can be represented by the basis vectors ${h_{0}^{(n)}}(\ket{\Psi_{0}}),\ldots,{h_{n-1}^{(n)}}(\ket{\Psi_{0}})$ of $\mathrm{H}_{n}^{n}$, it follows that $\mathrm{H}_{n}^{n}=\mathrm{K}_{n}$ holds true. ∎
Theorem 3 (Approximation of Krylov Space).
Assume everything like in Theo. 2. Then for a large $N_{L}\in\mathbb{N}$
$G_{N_{L}}=\mathrm{Span}(\ket{g_{0}},..,\ket{g_{{N_{L}}-1}})$ (28)
, with some |$\ket{g_{i}}=\exp(-iHt_{i})\ket{\Psi_{0}}$ it holds that
$\mathrm{H}_{N_{L}}^{N_{L}}\approx\mathrm{G}_{N_{L}}^{N_{L}}$ (29)
From this it follows then that $\mathrm{G}_{N_{L}}\approx\mathrm{K}_{N_{L}}$ holds as well and finally $\mathrm{G}_{m}\approx\mathrm{K}_{m}$.
Proof.
We consider a $N_{L}\gg m$. This leads to
$\mathrm{K}_{m}$ $=\mathrm{K}_{N_{L}}=\mathrm{H}_{N_{L}}^{N_{L}}$
$=\mathrm{Span}\Big{(}\sum_{j=0}^{{N_{L}}-1}{f^{j}(\ket{\Psi_{0}})}\frac{t_{0}^{j}}{j!},\ldots,\sum_{j=0}^{{N_{L}}-1}{f^{j}(\ket{\Psi_{0}})}\frac{t_{{N_{L}}-1}^{j}}{j!}\Big{)}$ (30)
For any $t>0$ and sufficiently large $N_{L}$ the series terms $t^{N_{L}}/N_{L}!\rightarrow 0$, such that the approximation
$\sum_{j=0}^{{N_{L}}-1}{f^{j}(\ket{\Psi_{0}})}\frac{t^{j}}{j!}\approx\sum_{j=0}^{\infty}{f^{j}(\ket{\Psi_{0}})}\frac{t^{j}}{j!}$ (31)
can be used. Inserting this into Eq. 30 results in
$H_{N_{L}}^{N_{L}}$ $=\mathrm{Span}\Big{(}\sum_{j=0}^{{N_{L}}-1}{f^{j}(\ket{\Psi_{0}})}\frac{t_{0}^{j}}{j!},\ldots,\sum_{j=0}^{{N_{L}}-1}{f^{j}(\ket{\Psi_{0}})}\frac{t_{{N_{L}}-1}^{j}}{j!}\Big{)}$
$\approx\mathrm{Span}\Big{(}\sum_{j=0}^{\infty}{f^{j}(\ket{\Psi_{0}})}\frac{t_{0}^{j}}{j!},\ldots,\sum_{j=0}^{\infty}{f^{j}(\ket{\Psi_{0}})}\frac{t_{{N_{L}}-1}^{j}}{j!}\Big{)}$
$\Leftrightarrow$ $K_{N_{L}}=H_{N_{L}}^{N_{L}}\approx G_{N_{L}}.$ (32)
[…]
$\mathrm{G}_{m}=\mathrm{Span}\Big{(}\ket{g_{i_{0}}},\ket{g_{i_{1}}},\ldots,\ket{g_{i_{m-1}}}\Big{)}\approx\mathrm{K}_{m}.$ (33)
The property of $\mathrm{H}_{N_{L}}^{N_{L}}=\mathrm{K}_{N_{L}}$ was shown in the prior example for any ${N_{L}}\in\mathbb{N}$. In the case of ${N_{L}}=\infty$ the approximation would reduce to $\mathrm{K}_{N_{L}}=\mathrm{G}_{N_{L}}$, which implies $\mathrm{G}_{m}=\mathrm{K}_{m}$. In our simulations we observed not only the approximation $\mathrm{G}_{m}\approx\mathrm{K}_{m}$ but the equality of these two spaces, i.e. $\mathrm{G}_{m}=\mathrm{K}_{m}$. The benefit of using the vectors |$\ket{g_{i}}$ as a basis instead of the vectors ${f^{i}(\ket{\Psi_{0}})}$ is that these vectors are normalized as |$|\ket{g_{i}}|=|\exp(-iHt_{i})\ket{\Psi_{0}}|=|\ket{\Psi_{0}}|=1$. |$\ket{g_{i}}$ are therefor robust to numerical rounding when checking linear independence. Additionally, with the exception of a global phase |$\ket{g_{i}}\rightarrow\exp(i\alpha)\ket{g_{i}}$, the vectors |$\ket{g_{i}}$ can be reconstructed through measurements.
Later, we will perform the Gram-Schmidt algorithm on $\mathrm{G}_{m}$ to obtain the orthonormal basis {|$\big{\{}\ket{w_{i}}\big{\}}_{i}$, such that
$\mathrm{G}_{m}$ $=\mathrm{Span}\Big{(}\ket{g_{0}},\ldots,\ket{g_{{m-1}}}\Big{)}=\mathrm{Span}\Big{(}\ket{w_{0}},\ldots,\ket{w_{{m-1}}}\Big{)}$ (34)
holds.
The following two lemmas will add to Theo. 2 to show that the basis can be computed through quantum mechanical measurements.
Lemma 1.
The vector space $\mathrm{G}_{m}$ is invariant under a phase $\alpha\in\mathbb{R}$.
Proof.
$e^{i\alpha}\mathrm{K}_{m}$ $=e^{i\alpha}\text{Span}(\ket{g_{0}},..,\ket{g_{m-1}})$
$=\text{Span}(e^{i\alpha}\ket{g_{0}},..,e^{i\alpha}\ket{g_{m-1}})=\mathrm{K}_{m}$ (35)
In the last step, we have used that $z=e^{i\alpha}\neq 0$ is a complex number. This can be extended to different complex phases $\alpha_{a}$.
$\mathrm{K}_{m}$ $=\text{Span}(e^{i\alpha_{0}}\ket{g_{0}},..,e^{i\alpha_{m-1}}\ket{g_{m-1}})=\mathrm{K}_{m}$ (36)
Lemma 2.
Consider a quantum system with Hamiltonian $H$ having eigenvectors |$\ket{\phi_{i}}$ and corresponding eigenvalues $\epsilon_{i}$ given by
$H\ket{\phi_{i}}=\epsilon_{i}\ket{\phi_{i}}.$ (37)
Let |$\ket{\Psi_{0}}$ be a non-trivial initial state, specifically a superposition of at least two eigenstates |$\ket{\phi_{n}}$ and |$\ket{\phi_{m}}$ with distinct eigenvalues $\epsilon_{n}\neq\epsilon_{m}$. Additionally, assume two distinct times $t_{1}$ and $t_{2}$, both less than the system’s period $T_{P}$. Then, it follows that the two time-evolved states |$\ket{g_{1}}=\exp(-iHt_{1})\ket{\Psi_{0}}$ and |$\ket{g_{2}}=\exp(-iHt_{2})\ket{\Psi_{0}}$ are linearly independent.
Proof.
The two states |$\ket{g_{1}}$ and |$\ket{g_{2}}$ can be represented as
|$\ket{g_{1}}$ $=e^{-iHt_{1}}\ket{\Psi_{0}}=\sum_{i}e^{-i\epsilon_{i}t_{1}}\ket{\phi_{i}}\bra{\phi_{i}}\ket{\Psi_{0}}$ (38)
|$\ket{g_{2}}$ $=e^{-iHt_{2}}\ket{\Psi_{0}}=\sum_{i}e^{-i\epsilon_{i}t_{2}}\ket{\phi_{i}}\bra{\phi_{i}}\ket{\Psi_{0}}.$ (39)
Substituting $\alpha_{i}=\bra{\phi_{i}}\ket{\Psi_{0}}$ leads to
|$\ket{g_{1}}$ $=e^{-iHt_{1}}\ket{\Psi_{0}}=\sum_{i}\alpha_{i}e^{-i\epsilon_{i}t_{1}}\ket{\phi_{i}}=(\ket{\phi_{1}},\ket{\phi_{2}},..,\ket{\phi_{N}})\begin{pmatrix}\alpha_{1}e^{i\epsilon_{1}t_{1}}\\ \alpha_{2}e^{i\epsilon_{2}t_{1}}\\ :\\ \alpha_{N}e^{i\epsilon_{N}t_{1}}\end{pmatrix}$ (40)
|$\ket{g_{2}}$ $=e^{-iHt_{2}}\ket{\Psi_{0}}=\sum_{i}\alpha_{i}e^{-i\epsilon_{i}t_{2}}\ket{\phi_{i}}=(\ket{\phi_{1}},\ket{\phi_{2}},..,\ket{\phi_{N}})\begin{pmatrix}\alpha_{1}e^{i\epsilon_{1}t_{2}}\\ \alpha_{2}e^{i\epsilon_{2}t_{2}}\\ :\\ \alpha_{N}e^{i\epsilon_{N}t_{2}}\end{pmatrix}$ (41)
The matrix representation $\mathbf{a}_{i}$ of the vector |$\ket{g_{i}}$ in regards to the eigenstates {|$\Big{\{}\ket{\phi_{i}}\Big{\}}$ is given by
$\mathbf{a}_{i}:=\begin{pmatrix}\alpha_{1}e^{i\epsilon_{1}t_{i}}\\ \alpha_{2}e^{i\epsilon_{2}t_{i}}\\ :\\ \alpha_{N}e^{i\epsilon_{N}t_{i}}\end{pmatrix}.$ (43)
Given the condition that there exist at least two distinct eigenvalues $\epsilon_{n}\neq\epsilon_{m}$ with non-zero coefficients $\alpha_{n}$ and $\alpha_{m}$ implies that the basis representations $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$ are independent. Due to this independence and the orthonormality of the eigenstates, it follows that the states |$\ket{g_{1}}$ and |$\ket{g_{2}}$ are linearly independent. Since any two states |$\ket{g_{i}}$ and |$\ket{g_{j}}$ are linearly independent for any distinct $t_{i}\neq t_{j}$, and as there exist a total of $m$ linearly independent vectors |$\ket{g_{i}}$ (as described in Theo. 2), it follows that for any $t_{0}<t_{1}<t_{2}<\ldots<t_{m-1}<T_{P}$, $m$ linearly independent vectors are obtained, represented by |$\ket{g_{i}}=\exp(-iHt_{i})\ket{\Psi_{0}}$. In the case of an equidistant time axis, meaning $t_{j}=(j+1)\tau/N$ for some $\tau<T_{P}$ and $i<m$, and with |$\ket{g_{j}}=\exp(-iHt_{j})\ket{\Psi_{0}}$, this results in
$\mathrm{G}_{m}=\text{Span}\Big{(}\ket{g_{0}},\ket{g_{1}},..,\ket{g_{m-1}}\Big{)}\approx\mathrm{K}_{m}.$ (44)
The space $\mathrm{G}_{m}$ comprises of states with the common initial state |$\ket{\Psi_{0}}$ and varying evolution times $t_{i}$. Notably, because of the invariance under a global phase (Lemma 1) $\mathrm{G}_{m}$ can be computed through quantum measurements.
Theo. 4 will discuss the similarities between $\mathrm{G_{m}}$ and a basis representation through the eigenstates of the system. We show that through a combination of the eigenstates of the Hamiltonian the time-evolution can be represented through $d$ superpositions of eigenstates of the system, where $d$ is the number of pairwise distinct eigenvalues. This will help us to understand the Krylov space and its meaning. If both spaces have the minimum number of basis states, then the dimension of these two spaces should be equivalent. This will further be numerically verified in Section 4.
Theorem 4.
Consider a Hamiltonian $H\in\mathbb{C}^{N\times N}$ with $H=H^{\dagger}$, possessing $d$ pairwise distinct eigenvalues, and satisfying the eigenvalue equation
$H\ket{\phi_{j}}=\epsilon_{j}\ket{\phi_{j}}.$ (45)
Then, it follows that the dimension (grade) of the Krylov space $\mathrm{K}_{m}$ is equivalent to the count of pairwise distinct eigenvalues $\epsilon_{j}$.
Proof.
With $\alpha_{j}=\bra{\phi_{j}}\ket{\Psi_{0}}$, it follows:
[…]
Since $H$ is an $N\times N$ matrix, there exist $N$ eigenvalues, where only $d$ are pairwise distinct. The $d$ pairwise distinct eigenvalues are $\epsilon_{1},\ldots,\epsilon_{d}$. As $H$ is Hermitian, it follows that the eigenstates |$\ket{\phi_{j}}$ form an orthogonal basis of $\mathbb{C}^{N}$. Assume that there are $j_{i}$ eigenvectors with the same eigenvalue $\epsilon_{i}$. Introduce the index sets $J_{1}=\{1,\ldots,j_{1}\}$ and $J_{k}=\{j_{k-1}+1,\ldots,j_{k}\}$ with $k\in\{2,\ldots,d\}$. The index sets group eigenstates |$\ket{\phi_{i}}$, so that all eigenstates |$\ket{\phi_{a}}$, with $a\in J_{k}$, have the same corresponding eigenvalue $\epsilon_{k}$. This can be used to rewrite the time evolution as:
|$\ket{\Phi(t)}$ $=\sum_{j}e^{-i\epsilon_{j}t}\alpha_{j}\ket{\phi_{j}}=\sum_{p=1}^{d}\sum_{j\in J_{p}}e^{-i\epsilon_{p}t}\alpha_{j}\ket{\phi_{j}}=\sum_{p=1}^{d}e^{-i\epsilon_{p}t}\sum_{j\in J_{p}}\alpha_{j}\ket{\phi_{j}}$ (47)
The summation over the index sets $J_{p}$ is independent of time $t$ and can therefore be substituted with vectors |$\ket{\xi_{p}}=\sum_{j\in J_{p}}\ket{\Phi_{i}}$, which results in
|$\ket{\Phi(t)}=\sum_{p=1}^{d}e^{-i\epsilon_{p}t}\ket{\xi_{p}}.$ (48)
Such that for all $t\in\mathbb{R}$ the following holds:
|$\ket{\Phi(t)}\in\text{Span}\Big{(}\xi_{1},\xi_{2},\ldots,\xi_{d}\Big{)}:=X_{d}.$ (49)
The space $X_{d}$ consists of a superposition of $d$ linearly independent vectors |$\ket{\xi_{1}},..,\ket{\xi_{d}}$. For any time any two vectors are linearly independent, which implies that $X_{d}$ is yet another minimal basis of the time evolution. Let $\mathrm{K}_{m}$ be the Krylov space for $H$ initiated with the state |$\ket{\Psi_{0}}$, where $m$ denotes its grade and is therefor a minimal basis. Both spaces $X_{d}$ and $\mathrm{K}_{m}$ consist of the smallest number of vectors needed to represent any time-evolved state |$\ket{\Phi(t)}$ with a initial condition |$\ket{\Psi_{0}}$ in regards to the Hamiltonian $H$. Therefore, it follows that $d=m$ must hold true. ∎
Up until now, we have shown in Theo. 3 that time-evolved states exist, which span the same space as $\mathrm{K}_{m}$. We further showed the phase invariance (Lemma 1) and that the dimension is equal to the number of pairwise distinct eigenvalues(Theo. 4). The next chapter will verify these results for simple systems.
4 Numerical verification
To check whether $\mathrm{G}_{m}$ fully captures the evolution of a state under the Schrödinger equation, we will check if any time-evolved state can be reconstructed by the basis of $\mathrm{G}_{m}$ without error. Furthermore, we want to make sure that the dimension of $\mathrm{G}_{m}$ is equal to the number of pairwise distinct eigenvalues $d$.
First, consider a Hamiltonian $H\in\mathbb{C}^{N\times N}$ along with $N_{S}$ randomly chosen initial states |$\ket{\Psi_{1}},\ket{\Psi_{2}},...,\ket{\Psi_{N_{S}}}$. Select a timescale $T<T_{P}$, where $T_{P}$ represents the period, where $\exp(-iH(T_{P}+t))=\exp(-iHt)$ holds. Proceed with a discretization $0=t_{0}<t_{1}<t_{2}<...<t_{N-1}<T$ and compute the vectors |$\ket{g_{i}}=\exp(-iHt_{i})\ket{\Psi_{x}}$ for a given starting state |$\ket{\Psi_{x}}$, where $x=1,...,N_{S}$. The matrix $A_{m}=(\ket{g_{0}},\ket{g_{1}},...,\ket{g_{m-1}})$ is constructed iteratively, incrementing $k$ until $rank(A_{m-1})=rank(A_{m})$ is reached for some $k=m$. Here, $m$ denotes the dimension of the vector space $\mathrm{K}_{m}$ and, by approximation, the dimension of $\mathrm{K}_{m}(-iH,\ket{\Psi_{x}})$ or the grade of |$\ket{\Psi_{x}}$ in regards to $f:=-iH$. After forming $\mathrm{G}_{m}=Span\Big{(}\ket{g_{0}},\ket{g_{1}},...,\ket{g_{m-1}}\Big{)}$, the Gram-Schmidt algorithm is applied to obtain orthonormalized vectors |$\ket{w_{0}},\ket{w_{1}},...,\ket{w_{m-1}}$, which results in
$\mathrm{G}_{m}$ $=Span\Big{(}\ket{g_{0}},...,\ket{g_{m-1}}\Big{)}=Span\Big{(}\ket{w_{0}},...,\ket{w_{m-1}}\Big{)}.$ (50)
Once the vectors |$\ket{w_{i}}$ are computed, $N_{T}$ random times $\tau_{1},\tau_{2},...,\tau_{N_{T}}$ are chosen, and states evolved by the times |$\ket{\Psi(\tau_{k})}=\exp(-iH\tau_{k})\ket{\Psi_{x}}$ are calculated. As per Theo. 1 |$\ket{\Psi(\tau)}\in\mathrm{K}_{m}$ is true for all $\tau\in\mathbb{R}$. Furthermore, Theo. 2 implies that $\mathrm{G}_{m}\approx\mathrm{K}_{m}$ is a valid approximation, which will be verified by demonstrating that for any time $\tau_{k}$, |$\ket{\Psi(\tau_{k})}\in\mathrm{G}_{m}$ holds true. The $l$-th approximation |$\ket{u_{l}(\tau_{k})}$ of |$\ket{\Psi(\tau_{k})}$ concerning the starting state |$\ket{\Psi_{x}}$ is given by
|$\ket{u_{l}(\tau_{k})}=\sum_{j=0}^{l-1}\bra{w_{j}}\ket{\Psi(\tau_{k})}\ket{w_{j}}.$ (51)
The error in the $l$-th approximation of a time-evolved state |$\ket{\Psi(\tau_{k})}$ with the initial vector |$\ket{\Psi_{x}}$ is quantified by
$r(\tau_{k},l,\ket{\Psi_{x}})=\Big{|}\ket{u_{l}(\tau_{k})}-\ket{\Psi(\tau_{k})}\Big{|}.$ | . (52)
If $\mathrm{G}_{m}=\mathrm{K}_{m}$ holds true, then $r(\tau_{k},m,\ket{\Psi_{x}})=0$ is expected. For a statistical analysis, multiple times $\tau_{k}$ are considered, and the average over these times is calculated as $r(l,\ket{\Psi_{x}})$.
$r(l,\ket{\Psi_{x}})=\frac{1}{N_{T}}\sum_{k=1}^{N_{T}}r(\tau_{k},l,\ket{\Psi_{x}}).$ (53)
Moreover, $N_{S}$ initial conditions |$\ket{\Psi_{x}}$, where $x=1,...,N_{S}$, are provided. After averaging over all possible initial conditions, the error of the $l$-th approximation is expressed as
$r(l)=\sum_{x=1}^{N_{S}}r(l,\ket{\Psi_{x}}).$ (54)
If $\mathrm{G}_{m}=\mathrm{K}_{m}$ is confirmed, where $m$ is chosen such that $r(m)=0$ for all |$\ket{\Psi_{x}}$ and all $\tau_{k}$, then the approximation is deemed valid. The number of distinct eigenvalues $d$ is expected to be equal to $m$, i.e., $m=d$, for all Hamiltonians. In an $N_{Q}$-qubit system, the $N_{Q}$-qubit Pauli matrices $P_{i}\in\{X_{i},Y_{i},Z_{i}\}$ are constructed as
$P_{i}=\bigotimes_{j=1}^{i-1}I_{2}\otimes\sigma_{p}\otimes\bigotimes_{j=i+1}^{N_{Q}}I_{2},$ (55)
where $p\in\{\sigma_{x},\sigma_{y},\sigma_{z}\}$, and $I_{2}$ represents the Pauli matrices and the 2-dimensional identity.
In this study, we conduct an analysis of various four-qubit Hamiltonians. Initially, attention is given to simple Hamiltonians denoted as $H_{1},H_{2},H_{3}$, and $H_{4}$, which consist of rotation in the x-direction of the first, first two, first three, and all four qubits respectively.
$H_{k}=\sum_{i=1}^{k}0.5X_{k}$ (56)
Subsequently, the analysis expands to include Ising Hamiltonian $H_{I1},H_{I2}$ and $H_{I3}$ with different inter-spin couplings $J_{ij}$.
$H_{I}=\sum_{i=1,j>i}^{4}J_{ij}X_{i}X_{j}+0.5\sum_{i=0}^{4}Z_{i}$ (57)
Further information on the couplings $J_{ij}$ is listed in Appendix A. Our objective is to assess whether the error $r(m)=0$ and if $m=d$ holds. Table 1 shows the calculated grade $m$ of the Krylov space and the number of pairwise distinct eigenvalues for all Hamiltonian. We observe in each case the equivalence between the two values, which gives numerical verification of Theo. 4.
$H_{1}$ $H_{2}$ $H_{3}$ $H_{4}$ $H_{I1}$ $H_{I2}$ $H_{I3}$
Grade of $\mathrm{G}_{m}$, $m$ 2 3 4 5 9 16 15
#eigenvalues, $d$ 2 3 4 5 9 16 15
Table 1: The grades $m$ of the different quantum systems and the number of pairwise distinct eigenvalues $d$. For each system equality is observed.
Figure 1.a) illustrates the averaged error of the reconstruction error $r(l)$ of the simple Hamiltonians $H_{i}$. In each scenario, reconstruction is evident when $l=m$, as indicated by $r(m)=0$ for each Hamiltonian. By increasing the complexity of the Hamiltonian from $H_{1}$ to $H_{2}$ and others, we increase the number of pairwise distinct eigenvalues and the dimension of the Krylov space, as shown in Table 1. Next, we compute the Krylov spaces and calculate the errors for some realizations of the Ising Hamiltonian, denoted as $H_{I1}$, $H_{I2}$, and $H_{I3}$ and show the reconstruction error $r(l)$ in Figure 1.c). Even though the Hamiltonian $H_{I3}$ seems more complex compared to $H_{I2}$, the Krylov dimension and the number of pairwise distinct eigenvalues are smaller. Here as well, the correct reconstruction of all states by the calculated spaces $G_{m}$ can be observed.
Figure 1: The $l$th averaged reconstruction error in a) and c) for the simple Hamiltonians $H_{1}$ to $H_{4}$ and Ising Hamiltonians $H_{I1}$ to $H_{I3}$ and the effective Krylov dimension in b) and d) for the same Hamiltonians.
These numerical results emphasize that the defined space $\mathrm{G}_{m}$ forms a robust basis resilient to numerical errors, a quality that can be further verified through quantum-mechanical measurements.
Next, an effective space dimension will be constructed to describe the space $G_{m}$. For the effective dimension, we will look at states after they are evolved by a time $T$. We compute states |$\ket{g_{i}}=e^{-iH\theta_{i}}\ket{\Psi_{0}}$ with evolution times $\theta_{i}=(i+1)T/m$ where $i=1,..,m$, and $m$ is the Krylov space dimension of the corresponding system. By Lemma 2, we know that all |$\ket{g_{i}}$ are linearly independent in the mathematical sense. If two vectors are very close to each other, i.e., |$\ket{g_{i}}\approx\ket{g_{i+1}}$, we require an effective dimension to consider the second vector. To obtain an effective dimension, we first calculate
$\lambda_{i}=\absolutevalue{\bra{g_{i}}\ket{g_{i+1}}}^{2}.$ (58)
The effective dimension $m_{eff}$ for $\lambda=1/\sqrt{2}$ is given by
$m_{{eff}_{i}}$ =$=\begin{cases}1&\text{if }\lambda_{i}<\lambda\\ 1-\frac{1}{1-\lambda}\cdot(\lambda_{i}-\lambda)&\text{if }\lambda_{i}\geq\lambda\end{cases}$
$m_{eff}$ $=1+\sum_{i=1}^{m-1}m_{{eff}_{i}}$ (59)
The first vector increases the dimension by $1$. Afterwards we perform the interpolation between two states |$\ket{g_{i}}$. $\lambda_{i}=1$ implies that the two vectors are linearly dependent and therefore numerically dependent, resulting in $m_{{eff}_{i}}=0$. $\lambda_{i}<\lambda$ implies that the two vectors are different enough to consider full numerical independence, and the effective dimension is then increased by $m_{{eff}_{i}}=1$. For $\lambda<\lambda_{i}\leq 1$ the effective dimension is interpolated. The measure $m_{{eff}_{i}}$ is upper-bounded by the Krylov space dimension $m_{{eff}_{i}}\leq m$ and is shown for different systems in Figure 1.b) and Figure 1.d). In the first row, we see the effective dimension for the four Hamiltonians $H_{1},H_{2},H_{3}$, and $H_{4}$. The effective dimension of $H_{1}$ (orange line) starts at one and then increases to the number of pairwise distinct eigenvalues, indicated by an orange horizontal line. Interestingly, we observe a decrease in effective dimension for $T=4\pi$. This can be explained by the $2\pi$ periodicity of the time evolution under $H_{1}$ as $\exp(-i2\pi X_{1}/2)=\exp(-iX_{1}/2)=-\exp(-i4\pi X_{1}/2)$. However, because the dimension of the Krylov space is $m=2$, we sample two time-evolved states at time $\theta_{1}=T/2$ and $\theta_{2}=T$. For $\theta=4\pi$, both are multiples of $2\pi$, which leads to the two vectors being |$\ket{g_{1}}=-\ket{g_{2}}$. |$\ket{g_{2}}$ being a scalar multiple of |$\ket{g_{1}}$ implies linear dependence, which in turn reduces the effective dimension $m_{\text{eff}}$. $H_{2}$ shows a small effective dimension for small times and increases for larger $T$. The vectors of $G_{m}$ are sampled at times $\theta_{1}=T/3,\theta_{2}=2T/3$, and $\theta_{3}=T$. The evolved states are linearly dependent for $T=3\cdot 2\pi$, as can be seen by the reduced effective dimension of the red curve for $T=3\cdot 2\pi$. The effective dimension of $H_{3}$ and $H_{4}$ follows the same trend, where it starts at one and then increases towards the number of pairwise distinct eigenvalues, where it saturates. In the second row, the effective dimension of the Ising Hamiltonians can be seen. An increase for larger times and then a saturation towards the number of pairwise distinct eigenvalues can be observed.
5 Discussion and Conclusion
In this work, we discuss Krylov spaces for pure states, building upon prior research on operator and state complexity as explored in PAR19 ; BAL22 ; ALI23 . The basis vectors of the already researched vector space $\mathrm{K}$ exhibit unequal lengths, posing challenges for orthonormalization. Additionally, a substantial number of matrix multiplications are required to compute the powers $H^{i}$, and rounding errors accumulate with each power. Since $H$ is a Hermitian matrix, experimental computation of $H^{i}$ is currently unfeasible, requiring classical machine computation. While manageable for small systems, this becomes challenging for large quantum systems.
These challenges find resolution with the introduction of the space $\mathrm{G}$, where each basis vector consists of a state evolved over time under the corresponding Hamiltonian, ensuring normalization. We first show the equivalence between the Krylov space $\mathrm{K}$ and the state space $\mathrm{G}$, see Theo. 2. Orthonormalization of the basis of $\mathrm{G}$ becomes more straightforward as all basis vectors are of equal length. Furthermore, each vector can be computed through a single matrix exponential operation, denoted as |$\ket{g_{j}}=e^{-iHt_{j}}\ket{\Psi_{0}}$, ensuring that each vector exhibits similar rounding errors. Theo. 4 shows that the dimension of $\mathrm{K}$ and $\mathrm{G}$ should be $d$, where $d$ represents the number of pairwise distinct eigenvalues of the Hamiltonian $H$.
Finally, we demonstrate the invariance of the vector space $\mathrm{G}$ under a global phase, which facilitates the computation of basis vectors |$\ket{g_{j}}$ through quantum mechanical measurements. This property allows for the effective utilization of this expressivity measure in large quantum mechanical systems and aids in determining the number of pairwise distinct eigenvalues in a quantum mechanical experiment, as illustrated in Lemma 1. To assess the representation accuracy of $\mathrm{G}$, we simulate various Hamiltonians and compute the dimension $m$ of $\mathrm{G}$ along with the number of pairwise distinct eigenvalues $d$ of the Hamiltonian in Section 4. For all simulated systems, we observe state representation at $m=d$, as indicated by a vanishing reconstruction error $r(m=d)=0$. This demonstrates that the defined space $\mathrm{G}$ is the smallest basis capable of representing any time-evolved state. We further expand on the description of $\mathrm{G}$ to define an effective dimension that characterizes the expressivity of the system, which can be of interest in understanding different quantum systems and quantum machine learning protocols.
The defined spaces play a crucial role in describing quantum mechanical systems. It would be particularly interesting to explore how this basis compares in spread complexity regarding the established powers of the Hamiltonian. Furthermore, this opens a new research question and a search for different basis representations that minimize spread complexity BAL22 . The defined effective dimension, bounded by the number of pairwise distinct eigenvalues, makes it possible to compare different quantum systems. This is especially important in quantum reservoir computing and quantum machine learning, where expressivity measures hold significant importance. In quantum reservoir computing, our preliminary research has shown that the effective Krylov dimension follows a similar trend to the information processing capacity for larger evolution times DAM12 ; MAR23 . Using the effective Krylov dimension in quantum machine learning training would also be of great interest for variational quantum circuits SCH15a ; MAT21 ; MCC16 ; SCH21 . Another research interest is to discuss measurable operator spaces and the influence of Krylov complexity PAR19 .
Appendix A Hamiltonians
$H_{1}$ $=0.5X_{1}$
$H_{2}$ $=0.5(X_{1}+X_{2})$
$H_{3}$ $=0.5(X_{1}+X_{2}+X_{3})$
$H_{4}$ $=0.5(X_{1}+X_{2}+X_{3}+X_{4})$
$H_{I1}$ $=0.5\sum_{i=1,j>i}^{3}X_{i}X_{j}+0.5\sum_{i=0}^{3}Z_{i}$
$H_{I2}$ $=0.4X_{1}X_{2}+0.5\left(\sum_{j=3}^{4}X_{1}X_{j}+\sum_{i=2,j>i}^{4}X_{i}X_{j}+\sum_{i=1}^{4}Z_{i}\right)$
$H_{I3}$ $=0.35X_{1}X_{2}+0.4X_{1}X_{3}+0.45X_{1}X_{4}+0.6X_{2}X_{3}$
$\quad+0.55X_{2}X_{4}+0.5X_{3}X_{4}+0.5\sum_{i=1}^{4}Z_{i}$